EXPLAIN ORDERED PAIRS TO ME. SUCH AS THIS PROBLEM Y=5+3X
Answer by prince_abubu(198)

An ordered pair, usually written as (x,y) are two numbers that if you plug them into an equation that asks for x and y, the equation will be true.

Let's take your equation y = 5 + 3x. An ordered pair for this equation would be (1,8). Another one would be (4,17). How'd we get that!? Actually, you can pick ANY x you want. This x-value that you have control over is sometimes called the INDEPENDENT variable because you have the freedom to choose whatever value you want for it. Once you choose your x-value, you would then have NO CONTROL of what the y-value would be because the equation takes over (or practically "decides" what to do with what you put in). Since you don't get to pick the y-value, the y-value is sometimes refered to as the DEPENDENT variable (because its value depends on what x-value you put in).

We chose a 1 for our x-value. We can't help but have a y-value that's 8. (That's because 5 + 3(1) = 8 = 7). The ordered pair would then be (1,8). When we chose 4 for the x, we couldn't help but get a 17 for the y (because 5 + 3(4) = 17 = y), so the ordered pair would be (4,17).